# What’s the Difference? Supporting Users in Identifying the Effects of Prompt and Model Changes Through Token Patterns

### v1

@@ 7.1 Story Generation with Llama
LLMs increasingly shape narratives Chung et al. (2022); Lin et al. (2024) and thus also our opinions and values Moyer-Gusé (2008); Schneider-Mayerson et al. (2023); Green and Appel (2024). It is, therefore, crucial to identify unexpected changes in LLM outputs that could cause harmful biases or propagate stereotypes. Here, we first explore story generation with the prompt "Tell a short story about a day in the life of a doctor." where we add a last name to the prompt (either Smith or Li) and where we upgrade the LLM to a more recent version.
Adding the last names to the prompt, Spotlight identifies expected patterns like (Smith) and (Li). It also identifies several unexpected systematic differences, like (He, he) when adding Smith. Similarly, when moving from Llama 2 to Llama 3 LLM with the Li prompt, it identifies the pattern (She, she).
@@
The name and model changes cause further systematic differences relating, e.g., to the type of doctor (general, practitioner), the description and behavior of the doctor (brilliant); (warm smile); (connect, personal, level), the place of work (small town); (clinic) or the linguistic tense (was, had). See Appendix Table 12 for details.
@@ 7.2 Experiments on GPT Outputs
Our experiments with GPT-4o-mini show further applications for Spotlight, e.g., to better understand how LLMs adapt their texts to a target audience or to guide interpretability research on multilingual LLMs by identifying differences of prompts in different languages. We also identify further unexpected effects. For story generation, we show, e.g., how patterns can indicate differences in cultural artifacts, like (Maria) and (Ramirez, Ramirez,) related to a shift in the background of protagonists (which a closer inspection reveals to not match actual census data). Or how, for a story about farming, changing the location from Kansas to Kenya causes changes in crops – (golden, wheat) vs. (maize) and (beans.) – but also affects the content of the story, focusing, e.g., more on (family).
Upgrading from GPT-3.5-turbo to GPT-4o-mini causes several surprising changes. For the medical story, over 60% of the texts suddenly talk about (oatmeal, breakfast). More problematic is the change for the Kenyan farming story where upgrading the LLM reduces the name diversity drastically. We discuss these and more discoveries in detail in Appendix D with all texts and patterns in the Additional Material.
@@ 8 User Study
A core element of our Spotlight workflow is the human user who wants to understand the effects of their LLM changes. We ran a user study to evaluate how difficult it is for users to identify significant differences and if they can leverage token patterns. We asked participants () on the crowdworking platform Prolific to identify differences between groups of LLM-generated texts in two settings, each with a target difference. The settings were inspired by the demonstration studies: one on the Dr.Li/Dr.Smith prompt (target difference: gender) and one on farming in Kenya and Kansas (target: culturally relevant crops). Participants saw the texts with or without a token pattern; setting order and intervention being randomized.
Figure 2 shows that identifying the effects of prompt differences is indeed challenging for users; less than 60% of participants identified the gender difference and only 25% found the crop difference. Also, for the farming setting, where the crops were the only real systematic difference, 54% of participants nevertheless reported other differences. This confirms the difficulty for users to distinguish between actual systematic differences and random LLM output variations. Providing token patterns to the user boosts the success rate in finding the target difference highlighting the usefulness of our approach for supporting the prompt engineering process. More details and results in Appendix E.
@@ 9 Discussion & Conclusion
In contrast to existing approaches, Spotlight does not restrict LLM outputs to easy-to-count one-word answers or labels, but instead embraces long, generated texts. We highlight the advantage of this approach both on existing prompt data, as well as in several demonstration studies we developed. Spotlight is able to identify a variety of fine-grained – and sometimes surprising – effects of prompt and model changes, e.g., how adding a last name in the prompt can cause drastic shifts in gender distribution or how using a specific persona can suddenly cause task failures in rare but relevant cases. We aspire for our approach to become a useful addition to the existing evaluation toolbox for prompt engineering and model behavior research.
Our user study showed that the proposed token patterns indeed support the user in discovering significant changes, and our study also identified interesting future, human-centered research questions on how to best leverage these for the user. Future work could also explore how Spotlight, which provides direct links to the text data, could be combined with the expressiveness of LLMs to communicate insights to the user even better.
@@ Acknowledgments
We express our gratitude to the members of the MaiNLP lab for their constructive feedback, especially Felicia, Philipp, Rob and Silvia. We gratefully acknowledge that experiments involving API calls to GPT models were supported by a compute grant from OpenAI. BP is supported by European Research Council (ERC) Consolidator Grant DIALECT 101043235.
@@
Given the qualitative results from the benchmarks, we decided for the evaluation with real prompting set-ups, to perform in-depth analyses in line with the goal of our method to support nuanced analysis. While we cover a diverse set of settings, like semantically insignificant prompt changes, personas, gender biases and cultural impacts, many more settings could be explored. We miss a wider, but possibly more shallow, evaluation. We make our tools publicly available in the hope that future work will provide even more, realistic evidence.
We ran our user study on a crowdworker platform to obtain a higher number of participants. This required, however, a more restricted testing setting to obtain quantifiable results. While our user study provides general evidence, it is limited in its insights on prompt engineering processes of specific users. A more in-depth, user-centric evaluation is needed here in the future that also includes aspects about the user interface and that could possibly require alternative, user-centric evaluation methods, like probe studies and quantitative analyses.
@@ Appendix A Token Pattern Discovery Methods
Both Cortana Meeng and Knobbe (2011) and PySubgroup Lemmerich and Becker (2018) are subgroup discovery methods that discover interesting subgroups in the data. A subgroup is identified by a logical formula on the features of the data. In our case, we use binary features about the presence/absence of a token in a text. The token pattern (he,him) would thus be represented as . We use the separation into the two groups as basis for the interestingness metric. We use the authors’ implementations in both cases using negative r as metric and min-coverage for Cortana and StandardQF as metric and SimpleDFS as search algorithm for PySubgroup. Both methods require defining the maximum length of patterns for which we use the ground truth length in the benchmarks.
@@ Appendix B Experiments With Known Ground Truth
We list here the details of the creation and evaluation of the benchmarks beyond what was already introduced in the main paper (Section 5) to better ensure reproducibility. We also provide all data for the benchmarks in the Additional Material. Table 2 lists data statistics for the benchmarks. We release the data of the benchmarks under the open Creative Commons BY 4.0 license.
@@ B.1 Benchmark 1
The runtime for all methods was below one minute. The spent budget for GPT-4o was below 2$.
@@ B.2 Benchmark 2
| Method | Example Patterns |
| Premise | (forward, hearing, from, look), (finds, well, hope, email), (plz, wanna), (Dear, Professor) |
| c-tf-idf | (to), (you), (and), (your) |
| PySubgroup | (Name, !, and, .), (with, !, ., I), (I, and, ., I), (Name, !, Thank, .) |
| Cortana | (Dear, Available), (like, personalized), (personalized, this), (Sincerely, manner) |
@@ B.3 Benchmark 3
The (strict) F1 score requires an exact match between the identified and the ground truth patterns. For example, (Dear, professor) must be detected exactly; identifying only (Dear) or (Dear, professor, Smith) would not be considered correct.
In contrast, the soft F1 score is more lenient. It considers a pattern correctly identified as long as it contains any overlapping words with the ground truth. Under this criterion, both (Dear) and (Dear, professor, Smith) would be accepted as correct detections. Precision and recall are computed analogously.
@@ C.1 Effects of Semantically Insignificant Changes
The results are discussed in Section 6.1 of the main paper with a subset of the patterns in Table 8. We provide the full LLM outputs and discovered patterns in the Additional Material.
@@
| (True) | 194 | 14 | True. Non-binary individuals do not identify as […] |
| (False) | 296 | 447 | False. The statement is false. […] |
| (||, :) | 0 | 116 | False. || Explanation: The course of human history […] |
@@
| (statement, implies) | 52 | 119 | […] The statement is false because it implies […] |
@@
| (However, note) | 14 | 62 | […] However, it is important to note that […] |
| (I, think) | 5 | 83 | I think the statement is […] |
@@
| (False) | 447 | 89 | False. The statement is false. […] |
| (True) | 14 | 172 | True. Competition can be a powerful motivator […] |
| (||,:) | 116 | 0 | False. || Explanation: The course of human history […] |
@@
| (I, think) | 24 | 83 | I think the statement is […] |
@@ C.2 Effects of Personas
Our approach retrieves token patterns that describe very plausible systematic differences on the semantic level when comparing the LLM outputs. Some patterns directly mirror the persona descriptions (e.g. (conservative), (highly, open, individual)). More patterns are closely related to the personas, like the conservative persona exhibiting patterns like (traditional, roles), (established, norms), (order, stability) and (been, for, centuries). Some aspects are not directly related to the persona description but might be associated with conservative values, like an emphasis on (taking, responsibility). For the "open" persona, we retrieve similarly patterns that are very plausible but go beyond the surface level of the persona description, like the answers highlighting the (complex, multifaceted) nature of these questions, issues with (power, dynamics) and the idea of (dismantle, systems) of oppression. Interestingly, while one might expect both a "conservative" and an "open" persona to use the term "biological sex" in their argument, this is much more prominent for the later persona.
Like it was studied by the original authors of the prompt questions, the patterns also indicate shifts in agreement, like (False) and (true, statement). Of special note is the pattern (partially, true, some, also, may) which identifies a setting ("conservative" persona, "discrimination" prompt) where the LLM outputs have a tendency to deviate from the clear (and originally intended) true/false answer in ca. 20% of the cases. This highlights again the need for such more nuanced evaluations.
@@
Most surprising was the pattern (You, you, believe) which occurs in some cases for the "agreeable" persona. Analyzing this token pattern shows that the LLM outputs fail the instructions and do not answer about the persona (first or third person), but instead about the reader (second person, "[…] you believe that a person’s gender […]"). With just 8% of outputs of the "agreeable" persona containing this failure (and none for the "conservative"), this is a difference that is relevant but which could have been easily missed when just fully manually evaluating a (human readable) subset of the LLM outputs. This underscores again the benefit of combining automatic and manual analysis through our token pattern approach.
More details about the discussed token patterns can be found in Table 10 with all LLM outputs and patterns in the Additional Material.
@@
| (traditional, roles) | 7 | 181 | […] traditional understanding of gender roles […] |
| (established, norms) | 0 | 130 | […] maintaining established norms and values. |
| (order, stability) | 0 | 90 | […] a threat to social order and stability. […] |
| (been, for, centuries) | 0 | 41 | […] expectations have been established for centuries […] |
| (You, you, believe) | 41 | 0 | […] you believe that a person’s gender […] |
@@
| (False) | 295 | 120 | False. Explanation: The statement implies […] |
| (true, statement) | 35 | 166 | […] Therefore, I believe the statement is true. |
| (highly, open, individual) | 157 | 0 | As a highly open individual […] |
| (biological, sex, typically) | 123 | 15 | […] While biological sex is typically seen […] |
| (complex, multifaceted) | 400 | 153 | […] identity is a complex and multifaceted construct […] |
@@
| (Furthermore) | 0 | 62 | […] Furthermore, the idea of "freezing" […] |
| (Moreover) | 3 | 40 | […] Moreover, the concept of gender […] |
| (fluid) | 37 | 201 | […] gender identity can be fluid […] |
| (frozen) | 445 | 282 | […] gender identity is not frozen in time […] |
@@
| (False) | 500 | 346 | False. The statement is false because […] |
| (partially, true, some, also, may) | 2 | 95 | […] is partially true but also partially false […] |
@@
| (taking, responsibility) | 44 | 92 | […] are not taking personal responsibility […] |
| (dismantle, systems) | 74 | 36 | […] dismantle systems of oppression. […] |
@@
| (I) | 2 | 92 | I think the statement is false. […] |
| (power, dynamics) | 10 | 40 | […] issue that involves power dynamics […] |
@@
| (conservative) | 0 | 489 | True. As a conservative, […] |
| (information) | 33 | 105 | […] lack of education, limited access to information […] |
@@
| (reflects, reality) | 0 | 125 | […] statement accurately reflects the reality of […] |
| (while, some) | 203 | 38 | […] False. While some individuals may face […] |
@@
| (For, example) | 11 | 55 | […] For example, a person living […] |
| (For, instance) | 8 | 39 | […] For instance, they may face […] |

### v2

@@ 7.1 Story Generation with Llama
LLMs increasingly shape narratives Chung et al. (2022); Lin et al. (2024) and thus also our opinions and values Moyer-Gusé (2008); Schneider-Mayerson et al. (2023); Green and Appel (2024). It is, therefore, crucial to identify unexpected changes in LLM outputs that could cause harmful biases or propagate stereotypes. Here, we first explore story generation with the prompt “Tell a short story about a day in the life of a doctor." where we add a last name to the prompt (either “Smith” or “Li”) and where we upgrade the LLM to a more recent version.
Adding the last names to the prompt, Spotlight identifies expected patterns like Smith and Li. It also identifies several unexpected systematic differences, like He, he when adding Smith. Similarly, when moving from Llama 2 to Llama 3 LLM with the Li prompt, it identifies the pattern She, she.
@@
The name and model changes cause further systematic differences relating, e.g., to the type of doctor general, practitioner, the description and behavior of the doctor brilliant; warm, smile; connect, personal, level, the place of work small town; clinic or the linguistic tense was, had. See Appendix Table 12 for details.
@@ 7.2 Experiments on GPT Outputs
Our experiments with GPT-4o-mini, detailed in Appendix D, show further applications for Spotlight, e.g., to better understand how LLMs adapt their texts to a target audience or to guide interpretability research on multilingual LLMs by identifying differences of prompts in different languages.
We also identify further unexpected effects for story generation, e.g., how patterns can indicate differences in cultural artifacts, like Maria and Ramirez, Ramirez, related to a shift in the background of protagonists (which a closer inspection reveals to not match actual census data). Or how, for a story about farming, changing the location from Kansas to Kenya causes changes in crops, golden, wheat vs. maize and beans., but also affects the content of the story, focusing, e.g., more on family.
Upgrading from GPT-3.5-turbo to GPT-4o-mini causes several surprising changes. For the medical story, over 60% of the texts suddenly talk about oatmeal, breakfast. More problematic is the change for the Kenyan farming story where upgrading the LLM reduces the name diversity drastically. For the older model, the protagonist’s name was distributed via a mixture of medium frequency and rare names. For the modern model, the outputs are dominated by a single name, indicated by the pattern Juma in a majority of the stories. We discuss these and more discoveries in detail in Appendix D.
@@ 8 User Study
A core element of our Spotlight workflow is the human user who wants to understand the effects of their LLM changes. We ran a user study to evaluate how difficult it is for users to identify significant differences and if they can leverage token patterns. We asked participants () on the crowdworking platform Prolific to identify differences between two groups of LLM-generated texts. Each pair of text groups contained one target difference. Participants were asked to describe the differences in free-text form, with or without a token pattern for the target difference being shown to them. They were told there could be no, one, or multiple differences. We tested two text scenarios and randomized the order of the scenarios and the token pattern intervention. The full experimental protocol is given in Appendix E.
A pilot study () informed the phrasing of the instructions and verified that the tasks are solvable. There, we identified that participants might be overwhelmed if exposed to too much text, mirroring insights from Gero et al. (2024). We thus prompted the LLM to limit outputs to 30 words and showed users 10 texts per group. We expect that for a setting with longer texts and more differences, this task would be even more difficult to users. The two scenarios are: 1) description of Dr.Li/Dr.Smith with the target difference being the gender and 2) description of the harvest of a farm where the difference are culturally relevant crop terms for Kensas and Kenya ("corn" vs. "beans" and "maize"). We limited participation to the US and UK and speakers of English as first language to reduce language skill effects. 3 participants were filtered out for failing quality checks. The participants’ answers were manually annotated.
Additionally, we surveyed participants about their background and experience when performing the task. 47% of participants reported that they use ChatGPT or LLMs at least once or twice per month, and 42% that they do data analysis tasks at least once or twice per month.
Identifying the effects of prompt differences is indeed challenging for users: less than 60% of participants (without token pattern) identified the gender difference and only 25% found the crop difference. Also, for the farming setting, where the crops were the only real systematic difference, 54% of participants nevertheless reported other differences. This confirms the difficulty for users to distinguish between actual systematic differences and random LLM output variations. Providing token patterns to the user boosts the success rate in finding the target difference, especially for the more difficult farming case (Figure 3), with further study details in Appendix E.
@@ 9 Discussion & Conclusion
In contrast to existing approaches, Spotlight does not restrict LLM outputs to easy-to-count one-word answers or labels, but instead works on long, generated texts. We highlight the advantage of this approach both on existing prompt data, as well as in several demonstration studies we developed. Spotlight is able to identify a variety of fine-grained—and sometimes surprising—effects of prompt and model changes, e.g., how adding a last name in the prompt can cause drastic shifts in gender distribution or how using a specific persona can suddenly cause task failures in rare cases.
Our user study showed that the proposed token patterns indeed support the user in discovering significant changes, but more in-depth, human-centered studies are necessary to explore the effects of Spotlight on the full prompt engineering process. Future work could also explore how Spotlight, which provides direct links to the text data, could be combined with the expressiveness of LLMs to better communicate insights to the user.
Making Spotlight’s code and resources publicly available, we aspire for our approach to become a useful addition to the existing evaluation toolbox for prompt engineering and model behavior research.
@@ Acknowledgments
We express our gratitude to the anonymous reviewers, the members of the MaiNLP lab (especially Felicia, Philipp, Rob and Silvia), and Anna, Adwait, David, Dawei, Jana, Joongi and Merle for their constructive feedback. We gratefully acknowledge OpenAI’s compute grant for calls to the GPT API. BP is supported by European Research Council (ERC) Consolidator Grant DIALECT 101043235.
@@
We cover common and crucial aspects of NLP analyses in our benchmarks, namely scalability (group size), bias ratio, pattern length (n-gram length) and word frequency. Other effect areas, however, might also play a role in LLM evaluation, such as high-level topics, information density or sequence order. These are not yet covered by our benchmarks. Additionally, our benchmarks are based on inserting systematic differences through token sequence replacement. Some of the yet to study effect areas might not be representable through such replacement procedures and especially the data mining methods might also struggle with identifying such non-token-based differences.
Having already qualitative results from the benchmarks, we decided for the evaluation with real prompting set-ups, to evaluate individual use cases in-depth in line with the goal of our method to support nuanced analysis. While we cover a diverse set of settings, like semantically insignificant prompt changes, personas, gender biases and cultural impacts, many more settings could be explored. While we apply Spotlight to evaluate model changes, we do not use it to evaluate hyperparameter changes that affect LLM outputs, like, e.g., temperature. In summary, we miss a wider, but possibly more shallow, evaluation. We make our tools publicly available in the hope that future work will provide even more, realistic evidence.
We ran our user study on a crowdworker platform to obtain a higher number of participants. This required, however, a more restricted testing setting to obtain quantifiable results. While our user study provides general evidence on the difficulty for users to identify systematic differences and that token-patterns can help with this task, it is limited to only two scenarios and it does not cover the full prompt engineering processes of specific users. Our participant group also tended towards being non-AI-experts. A wider user group would be necessary to get a more complete picture for different user groups, like LLM-engineers or scientists using LLMs. Studying different text lengths could also identify relevant usage differences. A more in-depth, user-centric evaluation is needed here in the future that also includes aspects about the user interface and that could possibly require alternative, user-centric evaluation methods, like probe studies and quantitative analyses.
@@ Appendix A Token Pattern Discovery Methods
Both Cortana Meeng and Knobbe (2011) and PySubgroup Lemmerich and Becker (2018) are subgroup discovery methods that discover interesting subgroups in the data. A subgroup is identified by a logical formula on the features of the data. In our case, we use binary features about the presence/absence of a token in a text. The token pattern he,him would thus be represented as . We use the separation into the two groups as basis for the interestingness metric. We use the authors’ implementations in both cases using negative r as metric and min-coverage for Cortana and StandardQF as metric and SimpleDFS as search algorithm for PySubgroup. Both methods require defining the maximum length of patterns for which we use the ground truth length in the benchmarks.
@@ Appendix B Experiments With Known Ground Truth
We list here the details of the creation and evaluation of the benchmarks beyond what was already introduced in the main paper (Section 5) to better ensure reproducibility. We also provide all data for the benchmarks in the Supplementary Material. Table 3 lists data statistics for the benchmarks. We release the data of the benchmarks under the open Creative Commons BY 4.0 license.
@@ B.1 Benchmark 1
Results are given in Tables 1 and 4. The runtime for all methods was below one minute. The spent budget for GPT-4o was below 2$.
@@ B.3 Benchmark 3
The (strict) F1 score requires an exact match between the identified and the ground truth patterns. For example, Dear, professor must be detected exactly; identifying only Dear or Dear, professor, Smith would not be considered correct.
In contrast, the soft F1 score is more lenient. It considers a pattern correctly identified as long as it contains any overlapping words with the ground truth. Under this criterion, both Dear and Dear, professor, Smith would be accepted as correct detections. Precision and recall are computed analogously.
@@ C.1 Effects of Semantically Insignificant Changes
The results are discussed in Section 6.1 of the main paper with a subset of the patterns in Table 8. We provide the full LLM outputs and discovered patterns in the Supplementary Material.
@@
| True | 194 | 14 | True. Non-binary individuals do not identify as […] |
| False | 296 | 447 | False. The statement is false. […] |
| ||, : | 0 | 116 | False. || Explanation: The course of human history […] |
@@
| statement, implies | 52 | 119 | […] The statement is false because it implies […] |
@@
| However, note | 14 | 62 | […] However, it is important to note that […] |
| I, think | 5 | 83 | I think the statement is […] |
@@
| False | 447 | 89 | False. The statement is false. […] |
| True | 14 | 172 | True. Competition can be a powerful motivator […] |
| ||,: | 116 | 0 | False. || Explanation: The course of human history […] |
@@
| I, think | 24 | 83 | I think the statement is […] |
@@ C.2 Effects of Personas
Our approach retrieves token patterns that describe very plausible systematic differences on the semantic level when comparing the LLM outputs. Some patterns directly mirror the persona descriptions (e.g. conservative, highly, open, individual). More patterns are closely related to the personas, like the conservative persona exhibiting patterns like traditional, roles, established, norms, order, stability and been, for, centuries. Some aspects are not directly related to the persona description but might be associated with conservative values, like an emphasis on taking, responsibility. For the "open" persona, we retrieve similarly patterns that are very plausible but go beyond the surface level of the persona description, like the answers highlighting the complex, multifaceted nature of these questions, issues with power, dynamics and the idea of dismantle, systems of oppression. Interestingly, while one might expect both a "conservative" and an "open" persona to use the term "biological sex" in their argument, this is much more prominent for the later persona.
Like it was studied by the original authors of the prompt questions, the patterns also indicate shifts in agreement, like False and true, statement. Of special note is the pattern partially, true, some, also, may which identifies a setting ("conservative" persona, "discrimination" prompt) where the LLM outputs have a tendency to deviate from the clear (and originally intended) true/false answer in ca. 20% of the cases. This highlights again the need for such more nuanced evaluations.
@@
Most surprising was the pattern You, you, believe which occurs in some cases for the "agreeable" persona. Analyzing this token pattern shows that the LLM outputs fail the instructions and do not answer about the persona (first or third person), but instead about the reader (second person, "[…] you believe that a person’s gender […]"). With just 8% of outputs of the "agreeable" persona containing this failure (and none for the "conservative"), this is a difference that is relevant but which could have been easily missed when just fully manually evaluating a (human readable) subset of the LLM outputs. This underscores again the benefit of combining automatic and manual analysis through our token pattern approach.
More details about the discussed token patterns can be found in Table 10 with all LLM outputs and patterns in the Supplementary Material.
@@
| traditional, roles | 7 | 181 | […] traditional understanding of gender roles […] |
| established, norms | 0 | 130 | […] maintaining established norms and values. |
| order, stability | 0 | 90 | […] a threat to social order and stability. […] |
| been, for, centuries | 0 | 41 | […] expectations have been established for centuries […] |
| You, you, believe | 41 | 0 | […] you believe that a person’s gender […] |
@@
| False | 295 | 120 | False. Explanation: The statement implies […] |
| true, statement | 35 | 166 | […] Therefore, I believe the statement is true. |
| highly, open, individual | 157 | 0 | As a highly open individual […] |
| biological, sex, typically | 123 | 15 | […] While biological sex is typically seen […] |
| complex, multifaceted | 400 | 153 | […] identity is a complex and multifaceted construct […] |
@@
| Furthermore | 0 | 62 | […] Furthermore, the idea of "freezing" […] |
| Moreover | 3 | 40 | […] Moreover, the concept of gender […] |
| fluid | 37 | 201 | […] gender identity can be fluid […] |
| frozen | 445 | 282 | […] gender identity is not frozen in time […] |
@@
| False | 500 | 346 | False. The statement is false because […] |
| partially, true, some, also, may | 2 | 95 | […] is partially true but also partially false […] |
@@
| taking, responsibility | 44 | 92 | […] are not taking personal responsibility […] |
| dismantle, systems | 74 | 36 | […] dismantle systems of oppression. […] |
@@
| I | 2 | 92 | I think the statement is false. […] |
| power, dynamics | 10 | 40 | […] issue that involves power dynamics […] |
@@
| conservative | 0 | 489 | True. As a conservative, […] |
| information | 33 | 105 | […] lack of education, limited access to information […] |
@@
| reflects, reality | 0 | 125 | […] statement accurately reflects the reality of […] |
| while, some | 203 | 38 | […] False. While some individuals may face […] |
@@
| For, example | 11 | 55 | […] For example, a person living […] |
| For, instance | 8 | 39 | […] For instance, they may face […] |